OpenStudy (anonymous):

y + 2x = −1 3y − x = 4

OpenStudy (anonymous):

using the first equation we can conclude that y=-2x-1 now we substitute this into our second equation 3(-2x-1)-x=4 -6x-3-x=4 -7x=7 x=-1 now we solve for we by pluggin in -1 into our first equation y=(-2(-1))-1 y=2-1 y=1